### v1

@@ Abstract
Figure S.6 demonstrates the typical time scaling for a Qdyne measurement Schmitt et al. (2017); Staudenmaier et al. (2021). The linewidth $2\sigma$ of the signal decreases exactly with 1/$T_{\mathrm{tot}}$ (see Fig. S.6(a)), as expected from theory. Figure S.6(b) shows that the precision $\delta\nu_{0}$ in frequency estimation scales with $T_{\mathrm{tot}}^{-3/2}$. Finally, SNR increases according to the standard quantum limit as $T_{\mathrm{tot}}^{1/2}$, as shown in Fig. S.6(c).
@@
We determine the frequency sensitivity of CPDD in combination with Qdyne by multiplying the estimated uncertainty $\delta\nu_{0}$ with the adequate time scaling, $\eta_{\nu_{0}}=\delta\nu_{0}\,T_{\mathrm{tot}}^{3/2}$, where $T_{\mathrm{tot}}$ is the total measurement time. This is characteristic for Qdyne, as shown in Fig. S.6(b) and Schmitt et al. (2017), to obtain 32 mHz/Hz3/2. Analogously, we calculate the amplitude sensitivity $\eta_{B}=\frac{B_{0}}{\mathrm{SNR}}T_{\mathrm{tot}}^{-1/2}$. We take the weighted average of the results in (d) and obtain $\eta$ = (118 $\pm$ 8) nT/$\sqrt{Hz}$.
@@
Next, we note that formulas for the frequency and amplitude sensitivity for CPDD are the same as for pulsed dynamical decoupling, except for a difference in the signal attenuation factor $\alpha$ (see also Section S.3 and Ref. Degen et al. (2017)). One can obtain intuition by considering the photon shot noise limited sensitivity for detecting the amplitude of an oscillating magnetic field with an NV center, using standard pulsed DD sequences the sensitivity is given by Degen et al. (2017); Taylor et al. (2008); Barry et al. (2020)
$\eta(t)=\frac{2}{\gamma_{\text{NV}}\alpha C(t)}\frac{1}{\sqrt{N_{\text{ph}}t}},$ (S.27)
@@
where $\gamma_{\text{NV}}$ is the NV center gyromagnetic ratio, $t$ is the sensing time, $N_{\text{ph}}$ is the the average number of photons collected per measurement run, $C(t)=(a-b)\exp{\left[-\Gamma(t)\right]}$ is the expected normalized contrast with $\Gamma(t)$ characterizing the respective decay due to the NV spin decoherence, $a$ and $b$ denote the normalized signal, which corresponds to the bright and dark states of the NV electron spin, respectively. The attenuation factor for pulsed dynamical decoupling is $\alpha=2/\pi$ Degen et al. (2017), while CPDD has an attenuation factor of $\alpha=1/2$, similar to the standard continuous dynamical decoupling protocol (see Eq. 3 in the main text). The optimal sensing time is $t=T_{2}/2$ for the most typical noise spectra Degen et al. (2017).
@@
The attenuation factor for CPDD is the same as with continuous dynamical decoupling and is slightly worse than with pulsed DD. In the latter case, the effect can be compensated by a longer coherence time as the repetition rate of the spin flips can be higher with CPDD than with pulsed methods due to the negligible waiting time. CPDD also has higher fidelity in comparison to continuous decoupling, which results in improved sensitivity.
@@
The dynamic range of CPDD is similar to the one of dressed states methods, e.g., continuous dynamical decoupling, i.e., 0.1-100 MHz for electronic spins or superconducting qubits Degen et al. (2017) with the additional advantage of improved robustness. We note that the dynamic range can be expanded substantially by combining CPDD with other techniques, similar to the possibility of sensing GHz fields with double drive continuous decoupling Cai et al. (2012); Salhov et al. (2024). Finally, we note that the reported sensitivities can be improved substantially by using an NV center with improved coherence properties, better photon collection efficiency or ensembles of NV centers where the average detected photon number per measurement is much higher and the effect of photon shot noise is thus significantly smaller.

### v2

@@ Abstract
Figure S.6 demonstrates the typical time scaling for a Qdyne measurement [41, 61]. The linewidth $2\sigma$ of the signal decreases exactly with 1/$T_{\mathrm{tot}}$ (see Fig. S.6(a)), as expected from theory. Figure S.6(b) shows that the precision $\delta\nu_{0}$ in frequency estimation scales with $T_{\mathrm{tot}}^{-3/2}$. Finally, SNR increases according to the standard quantum limit as $T_{\mathrm{tot}}^{1/2}$, as shown in Fig. S.6(c).
@@
We determine the frequency sensitivity of CPDD in combination with Qdyne by multiplying the estimated uncertainty $\delta\nu_{0}$ with the adequate time scaling, $\eta_{\nu_{0}}=\delta\nu_{0}\,T_{\mathrm{tot}}^{3/2}$, where $T_{\mathrm{tot}}$ is the total measurement time. This is characteristic for Qdyne, as shown in Fig. S.6(b) and [41], to obtain 32 mHz/Hz3/2. Analogously, we calculate the amplitude sensitivity $\eta_{B}=\frac{B_{0}}{\mathrm{SNR}}T_{\mathrm{tot}}^{-1/2}$. We take the weighted average of the results in (d) and obtain $\eta$ = (118 $\pm$ 8) nT/$\sqrt{Hz}$.
@@
Next, we note that formulas for the frequency and amplitude sensitivity for CPDD are the same as for pulsed dynamical decoupling, except for a difference in the signal attenuation factor $\kappa$ (see also Section S.3 and Ref. [6]). One can obtain intuition by considering the photon shot noise limited sensitivity for detecting the amplitude of an oscillating magnetic field with an NV center, using standard pulsed DD sequences. The sensitivity is given by [6, 52, 66]
$\eta(t)=\frac{2}{\gamma_{\text{NV}}\kappa C(t)}\frac{\sqrt{t+t_{\text{oh}}}}{\sqrt{N_{\text{ph}}}t},$ (S.27)
@@
where $\gamma_{\text{NV}}$ is the NV center gyromagnetic ratio, $t$ is the sensing time, $t_{\text{oh}}$ is the overhead time for readout and initialization, $N_{\text{ph}}$ is the the average number of photons collected per measurement run, $C(t)=(a-b)\exp{\left[-\Gamma(t)\right]}$ is the expected normalized contrast with $\Gamma(t)$ characterizing the respective decay due to the NV spin decoherence, $a$ and $b$ denote the normalized signal, which corresponds to the bright and dark states of the NV electron spin, respectively. For the NV center of the Qdyne measurements we estimate exponential decay of the signal with $\exp{\left[-\Gamma(t)\right]},\,\Gamma(t)=t/T_{2}$ and coherence time $T_{2}\approx 250\,\mathrm{\mu s}$. The average number of fluorescence photons per readout is $N_{\text{ph}}=0.164$ and the normalized contrast $a-b=0.32$. The sensing time is $t=5\,\mathrm{\mu s}$ corresponding to order 10 of a CXY8 at 8 MHz resonance frequency, and the overhead time $t_{\text{oh}}=3.646\,\mathrm{\mu s}$. With $\gamma_{\text{NV}}=2\pi\times 28.03\,\mathrm{MHz/mT}$ and $\kappa=1/2$ the theoretical photon shot noise limited sensitivity according to Eq. (S.27) is $\eta\approx 105\,\mathrm{nT/\sqrt{Hz}}$. The experimental sensitivity is slightly worse due to additional noise, that is not accounted for in Eq. (S.27).
The photon shot noise limited amplitude sensitivity could be improved further by optimizing the sensing time $t$ to minimize the expression $\frac{\sqrt{t+t_{\text{oh}}}}{C(t)t}$. We obtain an optimal sensing time of $t=\frac{1}{2}\left(\frac{T_{2}}{2}-t_{\text{oh}}+\sqrt{\frac{T_{2}^{2}}{4}+3T_{2}t_{\text{oh}}+t_{\text{oh}}^{2}}\right)\approx\frac{T_{2}}{2}+t_{\text{oh}}-\frac{4t_{\text{oh}}^{2}}{T_{2}}\approx 128.45\,\mathrm{\mu s}$, where the first approximation is valid for $t_{\text{oh}}\ll T_{2}$, reaching the theoretical optimum of $\eta_{\text{opt}}=26\,\mathrm{nT/\sqrt{Hz}}$. Finally, we note that the reported sensitivities can be improved substantially by using NV centers with longer coherence times, which can reach up to several milliseconds [26]. Other possibilities include improving further the photon collection efficiency or using ensembles of NV centers where the average detected photon number per measurement is much higher and the effect of photon shot noise is reduced significantly [6, 66].
@@
The signal attenuation factor for CPDD is $\kappa=1/2$, similarly to the standard continuous dynamical decoupling protocol (see Eq. (3)) in the main text). In contrast, pulsed dynamical decoupling has a slightly better attenuation factor of $\kappa=2/\pi$ [6]. This effect can be compensated by a longer coherence time as the repetition rate of the spin flips can be higher with CPDD than with pulsed methods due to the negligible waiting time. CPDD also has higher fidelity in comparison to continuous decoupling, which results in improved sensitivity.
@@
The dynamic range of CPDD is similar to the one of dressed states methods, e.g., continuous dynamical decoupling, i.e., 0.1-100 MHz for electronic spins or superconducting qubits [6] with the additional advantage of improved robustness. We note that the dynamic range can be expanded substantially by combining CPDD with other techniques, similar to the possibility of sensing GHz fields with double drive continuous decoupling [23, 26].